10. If we are doing two-sample hypothesis testing with related samples of sizes n1 = 10 and n2 = 10, what is the number of degrees of freedom?
a. 9
b. 10
c. 19
d. 20

11. If we are doing two-sample hypothesis testing with independent samples of sizes n1 = 10 and n2 = 10, what is the number of degrees of freedom?
a. 10
b. 18
c. 19
d. 20
Answer #1

10.

n1 = 10 and n2 = 10

Degrees of freedom = n2-1

DF = 10-1 = 9

11.

n1 = 10 and n2 = 10

Degrees of freedom = n1+n2-2

DF = 10+10-2 = 18